# Order of Addition in Mixture-Amount Experiments

### v1

@@ Abstract
In a mixture experiment, we study the behavior and properties of $m$ mixture components, where the primary focus is on the proportions of the components that make up the mixture rather than the total amount. Mixture-amount experiments are specialized types of mixture experiments where both the proportions of the components in the mixture and the total amount of the mixture are of interest. Such experiments consider the total amount of the mixture as a variable. In this paper, we consider an Order-of-Addition (OofA) mixture-amount experiment in which the response depends on both the mixture amounts of components and their order of addition. Full Mixture OofA designs are constructed to maintain orthogonality between the mixture-amount model terms and the effects of the order of addition. These designs enable the estimation of mixture-component model parameters and the order-of-addition effects. The G-efficiency criterion assesses how well the design supports precise and unbiased estimation of the model parameters. The Fraction of Design Space (FDS) plot is used to provide a visual assessment of the prediction capabilities of a design across the entire design space.
Keywords: Mixture experiments; Order-of-addition; Mixture-Amount Experiment; G-efficiency; Simplex-Lattice and Simplex-Centroid designs.
@@ 1 Introduction
In product formulations, the combination of various ingredients is observed, which forms a mixture. For example, chocolate is made of many ingredients in different compositions, cement-producing companies use certain ingredients in different proportions for making cement, backing shops use different ingredients for baking a cake, etc. The main focus of a mixture experiment is to judge an optimal formulation of a mixture and to get an optimal response from a composition of certain ingredients. An experiment in which the response variable $y$ depends on the proportion $x_{i}$ of components constituting a mixture blend, which is called a mixture experiment with the constraints
$0\leq x_{i}\leq 1\hskip 5.69046pt\text{and}\hskip 5.69046pt\sum_{i=1}^{n}x_{i};i=1,\dots,m$
[…]
A mixture-amount experiment is generally performed at two or more levels of the total amount, say $A$. The response is assumed to depend on the individual proportions of components in the blend and the amount of blend. The effects of varying the mixture component proportions and the total amount of the mixture on the response are measured by fitting a mixture-amount model to the design.
@@
This study focuses on the order in which the mixture proportions $x_{1},x_{2},\dots,x_{m}$ or component amounts $a_{1},a_{2},\dots,a_{m}$ are added in each blend. We assume that the blends add the $m$ components individually. Consequently, there are $m!$ possible sequences of these proportions or amounts to investigate. This scenario is referred to as the order-of-addition (OofA) problem in a mixture-amount experiment. The response variable $y$ depends on the order in which the component’s proportions or amounts are added in a composition and the interaction effect between the order and component proportions or amounts. In general, there are many scientific applications where the order of addition of the component proportions or amounts in a mixture blend affects the response.
Many physical phenomena observed in science and engineering are influenced by the order in which $m$ components or materials are added. Perhaps the first famous OofA experiment was performed by Fisher (1971) that of tasting tea by a lady who said that she could differentiate whether tea or milk was first added to her cup. The experiment consisted of four replications of each tea order $\rightarrow$ milk and milk $\rightarrow$ tea. The OofA experiments have found wide application also in bio-chemistry (Shinohara and Ogawa, 1998), nutritional science (Karim et al., 2000), bio-engineering (Chandrasekaran et al., 2006), chemical experiments (Jiang and Ng, 2014), combinatorial drug therapy (Ding et al., 2015), and many others.
Limited literature exists on Order-of-Addition (OofA) mixture experiments. Rios and Lin (2022) explored OofA mixture experiments and derived designs that account for interactions between the order of addition and mixture proportions. However, no prior research has addressed OofA mixture-amount experiments or their design construction. In our study, we develop designs that vary both the order of addition and the total mixture amount. These designs enable us to observe response changes for specific permutations of addition orders and total mixture amounts.
@@
A mixture-amount experiment examines how the response of a system is influenced by the individual proportions of components in a blend and the overall quantity of that blend, playing a crucial role in various fields such as pharmaceuticals, agriculture, nutrition, chemical engineering, and bioengineering. The primary objective is to determine the optimal proportions of ingredients to achieve the desired response, such as tablet weight and efficacy in the formulation of a pain relief tablet composed of active pharmaceutical ingredients, binders, and disintegrants. This involves designing experiments using factorial or mixture designs, identifying key variables like component proportions and total mixture amount, and measuring responses through statistical analysis to uncover significant effects on outcomes. The insights gained from these experiments can enhance product formulation and optimization, leading to improved performance and effectiveness across diverse applications, from developing optimal drug formulations to creating balanced nutrient profiles in fertilizers and food products.
@@ 1.1 Models for Mixture-Amount Experiments
The design for fitting the mixture-amount model is called the mixture-amount design, developed by Piepel and Cornell (1985). The linear and quadratic expected response mixture-amount models for $m$ components is
$E(y)$ $=$ $\sum_{i=1}^{m}\gamma_{i}^{0}x_{i}+\sum_{i=1}^{m}\gamma_{i}^{1}x_{i}A$ (1)
$E(y)$ $=$ $\sum_{i=1}^{m}\gamma_{i}^{0}x_{i}+\sum_{i<j}^{m}\gamma_{ij}^{0}x_{i}x_{j}+\sum_{l=1}^{2}\left(\sum_{i=1}^{m}\gamma_{i}^{l}x_{i}+\sum_{i<j}^{m}\gamma_{ij}^{l}x_{i}x_{j}\right)A^{l}$ (2)
This model consists of three second-order Scheffè polynomial forms in $m$ components, each multiplied by the powers of the total amount $A$ ($A^{0}=1$, $A$ and $A^{2}$). When $A=0$, this model predict a zero response. Piepel (1988) modified this model to accommodate zero-amount condition. The alternative model uses the actual amounts of the components denoted by $a_{1},a_{2},\dots,a_{m}$ such that $a_{1}+a_{2}\dots+a_{m}=A$. The proportions $x_{i}$ are related to the amounts $a_{i}$ through $x_{i}=a_{i}/A$ such that $a_{i}=x_{i}A$. This is called the component-amount model. The linear and quadratic expected response component-amount models are
@@
For further discussion on mixture-amount and component-amount models, see, for instance, Cornell (2002, pp. 403-418)
@@ 1.2 Designs for Mixture-Amount Experiments
Another mixture design, as discussed in Cornell (2002), is the simplex-centroid design. For an $m$-ingredient mixture, the simplex-centroid design consists of all vertices of the $(m-1)$-dimensional simplex, $S^{m-1}$, and centroids formed by taking these vertices in combinations, such as pairs. This design includes $12^{m}-$ 1 design points, starting with $m$ permutations of $(1,0,\dots,0)$, permutations of $\left(\frac{1}{2},\frac{1}{2},\dots,0\right)$, permutations of $\left(\frac{1}{3},\frac{1}{3},\frac{1}{3},\dots,0\right)$, continuing until the centroid $\left(\frac{1}{m},\frac{1}{m},\dots,\frac{1}{m}\right)$ is reached. Prescott and Draper (2004, 2008) developed several component-amount designs by projecting standard symmetric mixture designs, achieved by removing one or more columns from the original design. Additionally, orthogonally blocked mixture designs based on Latin squares and F-squares blocking schemes were projected to create orthogonal blocked component-amount designs.
[…]
@@ 2 OofA Experiments
The current literature on OofA experiments focuses primarily on the Pair-Wise Ordering model, first introduced by (Van Nostrand, 1995). The Pair-Wise Ordering (PWO) model was formally named by Voelkel (2019). For a set of $m$ components, their permutation is denoted by $\mathbf{c}=(c_{1},c_{2},\dots,c_{m})^{T}$. Let $\mathcal{P}$ be the set of all pairs $(j,k)$ where $1\leq j<k\leq m$.The PWO factor $z_{jk}(\mathbf{c})$ is defined as
$z_{jk}(\mathbf{c})=\begin{cases}1&\text{if $j$ precedes $k$ in $c$}\\ -1&\text{if $k$ precedes $j$ in $c$}\end{cases}$
@@
The parameter $\beta_{jk}$ shows how the pair-wise order of components $j$ and $k$ affect the response. For finding the optimal order Lin and Peng (2019) discussed the topological sorting methods for PWO model. There are several research papers on the Optimality of PWO experimental designs. See, for instance, Peng et al. (2019); Winker et al. (2020). Yuna Zhao and Liu (2021) and Chen et al. (2020).
@@ 2.1 The OofA mixture-amount experiments
An OofA mixture-amount experiment is a type of mixture experiment where we consider not only the effects of the component amounts on the response but also the effects of the order in which the components are added. So the response surface would be a function of components amounts $a_{1},a_{2},\dots,a_{m}$ and the order of their addition in a mixture. We test the significant effect of the order of addition of components and the amounts of components on the response. Our concern in this study is how to incorporate OofA effects into the mixture-amount and component-amount models. Therefore, we construct the OofA mixture component-amount designs using mixture simplex designs. We use the modified PWO notations given by Rios and Lin (2022). Define $(m-1)$ dimensional simplex $S$ and $\mathcal{P}$ be the set of all pairs $(j,k)$ where $1\leq j<k\leq m$. To get the full design matrix for fitting the mixture-amount model, we define PWO factor $z_{jk}(x,\mathbf{c})$ as
$z_{jk}(x,\mathbf{c})=\begin{cases}1&x_{j},x_{i}\neq 0\hskip 2.84544pt\text{$j$is before $k$ in $c$}\\ 0&x_{j}=0\hskip 2.84544ptor\hskip 2.84544ptx_{i}=0\\ -1&x_{j},x_{i}\neq 0\hskip 2.84544pt\text{$k$ is before $j$ in $c$}\\ \end{cases}$
@@
$z_{jk}(a,\mathbf{c})=\begin{cases}1&a_{j},a_{i}\neq 0\hskip 2.84544pt\text{$j$is before $k$ in $c$}\\ 0&a_{j}=0\hskip 2.84544ptor\hskip 2.84544pta_{k}=0\\ -1&a_{j},a_{i}\neq 0\hskip 2.84544pt\text{$k$ is before $j$ in $c$}\\ \end{cases}$
@@ 2.2 Models for OofA mixture-amount experiments
We further construct the model for the PWO and mixture component amounts. Without interaction between mixture-amount and PWO factor, the expected response additive model is developed as

### v2

@@ Abstract
In a mixture experiment, we study the behavior and properties of $m$ mixture components, where the primary focus is on the proportions of the components that make up the mixture rather than the total amount. Mixture-amount experiments are specialized types of mixture experiments where both the proportions of the components in the mixture and the total amount of the mixture are of interest. In this paper, we consider an Order-of-Addition (OofA) mixture-amount experiment in which the response depends on both the mixture amounts of components and their order of addition. Full mixture OofA designs are constructed to maintain orthogonality between the mixture-amount model terms and the effects of the order of addition. But the number of runs in such full OofA designs increases as $m$ increases. We employ the Threshold Accepting (TA) Algorithm to select an n-row subset from the full Order-of-Addition (OofA) mixture design that maximizes G-optimality while minimizing the number of experimental runs. Further, the G-efficiency criterion is used to assess how well the design supports the precise and unbiased estimation of the model parameters. These designs enable the estimation of mixture-component model parameters and the order-of-addition effects. The Fraction of Design Space (FDS) plot is used to provide a visual assessment of the prediction capabilities of a design across the entire design space.
keywords:
Mixture experiments, Order-of-addition, Mixture-Amount Experiment, G- efficiency, Simplex-Lattice and Simplex-Centroid designs
@@ 1 Introduction
The pharmaceutical industry relies heavily on research and development to produce effective and safe medications. Reducing costs and development time is crucial for maintaining competitiveness in this highly regulated and innovation-driven sector [1]. Medicines, which are carefully formulated mixtures of active pharmaceutical ingredients and excipients, must meet stringent quality standards to ensure therapeutic efficacy, stability, and patient safety. The optimization of these formulations requires a deep understanding of both the physicochemical properties of the components and their interactions.
A critical challenge in pharmaceutical research is determining the optimal proportions of mixture components to achieve desired therapeutic effects while minimizing adverse reactions. This problem is analogous to other industries where mixture optimization is essential, such as in baking (where ingredient ratios affect texture and taste) or construction (where material blends determine strength and durability) [2]. Even minor variations in component proportions in pharmaceuticals can significantly alter a drug’s dissolution rate, bioavailability, and overall effectiveness [3].
Mixture experiments provide a structured approach to optimizing formulations by systematically exploring how different component ratios influence key response variables, such as drug release kinetics or stability [4]. An experiment in which the response variable $y$ depends on the proportion $x_{i}$ of components constituting a mixture blend, which is called a mixture experiment with the constraints
$0\leq x_{i}\leq 1\hskip 5.69046pt\text{and}\hskip 5.69046pt\sum_{i=1}^{m}x_{i};i=1,\dots,m.$
The factor space of the mixture experiment in $m$ components is a (m1)(m-1) dimensional simplex $S^{m-1}$. Factor spaces resulting from specified constraints on mixture components belong to the simplex $S^{m-1}$. The $n$ points in the simplex $S^{m-1}$ or its sub-region generate a mixture design with $n$ runs. These $n$ runs are used to fit the mixture model to the observed responses. The design and analysis of such experiments enable researchers to identify the most effective formulations with minimal experimental runs, thereby reducing costs and development time [5].
@@
A further extension of this approach is the mixture-amount experiment, where the component proportions and the total quantity of the mixture are varied to study their combined effects on the response. This is particularly relevant in drug development, where different dosages may be required for varying patient populations. For instance, the impact of changing the mixture component proportions and the total amount of the mixture on the response is measured by fitting a mixture-amount model to the design. This study focuses on the order in which the mixture proportions $x_{1},x_{2},\dots,x_{m}$ or each component amount $a_{1},a_{2},\dots,a_{m}$ are added in each blend. We assume that the blends individually add the $m$ components. Consequently, there are $m!$ possible sequences of these proportions or amounts to investigate. This scenario is referred to as the order-of-addition (OofA) problem in a mixture-amount experiment. The response variable $y$ depends on the order in which the component’s proportions or amounts are added in a composition and the interaction effect between the order and component proportions or amounts. In general, there are many scientific applications where the order of addition of the component proportions or amounts in a mixture blend affects the response. For instance, physical phenomena observed in science and engineering are influenced by the order in which $m$ components or materials are added. The first famous OofA experiment was conducted by [6], which involved a lady who claimed to be able to differentiate whether tea or milk was added first to her cup. The experiment consisted of four replications of each tea order $\rightarrow$ milk and milk $\rightarrow$ tea. On the other hand, OofA experiments have also found wide application in biochemistry [7], nutritional science [8], bioengineering [9], chemical experiments [10], combinatorial drug therapy [11], and many others.
@@
Limited literature exists on OofA mixture experiments. Rios and Lin [12] explored OofA mixture experiments and derived designs that account for interactions between the order of addition and mixture proportions. If a mixture design with $m$ components has $t$ runs, then the OofA mixture design constructed would have $t\times m!$ runs. The number of runs increases exponentially as $m$ increases. Rios et al. [13] used the Threshold Accepting (TA) algorithm to find a subset that has $t\times m!$ runs based on the D-optimality criterion. However, no prior research has addressed OofA mixture-amount experiments and their design construction. In our study, we construct designs where both the order-of-addition and the total mixture amount are varied. Such OofA mixture component-amount designs would have $t$ x $m!$ mixture or component- order combinations. This paper focuses on designing cost-effective mixture-amount OofA experiments when funding constraints limit the number of runs. It is achieved by using the TA algorithm to select a limited set of runs from $t$ x $m!$ mixture or component-order combinations based on the G-optimality criterion. These designs enable us to observe response changes for specific permutations of addition orders and total mixture amounts. The primary objective is determining the optimal proportions of ingredients to achieve the desired response, such as tablet weight and efficacy in formulating a pain relief tablet composed of active pharmaceutical ingredients, binders, and disintegrants. This involves designing experiments using factorial or mixture designs, identifying key variables like component proportions and total mixture amount, and measuring responses through statistical analysis to uncover significant effects on outcomes. The insights gained from these experiments can enhance product formulation and optimization, leading to improved performance and effectiveness across diverse applications, from developing optimal drug formulations to creating balanced nutrient profiles in fertilizers and food products.
@@ 1.1 Models for Mixture-Amount Experiments
The design for fitting the mixture-amount model is called the mixture-amount design, developed by [14]. The linear and quadratic expected response mixture-amount models for $m$ components are
$E(y)=\sum_{i=1}^{m}\gamma_{i}^{0}x_{i}+\sum_{i=1}^{m}\gamma_{i}^{1}x_{i}A$
$E(y)$ $=$ $\sum_{i=1}^{m}\gamma_{i}^{0}x_{i}+\sum_{i<j}^{m}\gamma_{ij}^{0}x_{i}x_{j}+\sum_{l=1}^{2}\left(\sum_{i=1}^{m}\gamma_{i}^{l}x_{i}+\sum_{i<j}^{m}\gamma_{ij}^{l}x_{i}x_{j}\right)A^{l}.$ (1)
This model consists of three second-order Scheffè polynomial forms in $m$ components, each multiplied by the powers of the total amount $A$ ($^{0}=1, AA$ and $A^{2}$). When $A=0$, this model predicts a zero response. Piepel [15] modified this model to accommodate zero-amount condition. The alternative model uses the actual amounts of the components denoted by $a_{1},a_{2},\dots,a_{m}$ such that $a_{1}+a_{2}\dots+a_{m}=A$. The proportions $x_{i}$ are related to the amounts $a_{i}$ through $a_{i}=x_{i}A$. This is called the component-amount model. The linear and quadratic expected response component-amount models are
@@
For further discussion on mixture-amount and component-amount models, see, for instance, Cornell [2, pp. 403-418]
@@ 1.2 Designs for Mixture-Amount Experiments
Another mixture design, as discussed in [2], is the simplex-centroid design. For an $m$-ingredient mixture, the simplex-centroid design consists of all vertices of the (m1)(m-1)-dimensional simplex, $S^{m-1}$, and centroids formed by taking these vertices in combinations, such as pairs. This design includes $2^{m}-1$ design points, starting with $m$ permutations of $(1,0,\dots,0)$, permutations of $\left(\frac{1}{2},\frac{1}{2},\dots,0\right)$, permutations of $\left(\frac{1}{3},\frac{1}{3},\frac{1}{3},\dots,0\right)$, continuing until the centroid $\left(\frac{1}{m},\frac{1}{m},\dots,\frac{1}{m}\right)$ is reached. Prescott and Draper [16, 17] developed several component-amount designs by projecting standard symmetric mixture designs, achieved by removing one or more columns from the original design. Additionally, orthogonally blocked mixture designs based on Latin squares and F-squares blocking schemes were projected to create orthogonal blocked component-amount designs.
Any component-amount design with $m$ components can be derived from a simplex-lattice, simplex-centroid, or other standard mixture designs with $m+r$ ingredients by removing $r$ columns from the base design. For example, in a $\{4,3\}$ simplex-lattice design, deleting one column ($r=1$) produces a component-amount design with three ingredients and four levels of total amount $A$ (0, $1/3$, $2/3$, 1), denoted $\{4,3\}_{1}$. Similarly, we can derive a component-amount design with projection to three dimensions (deleting any column) from a simplex-centroid design with $q=4$ ingredients. The design has five levels of total amount AA (AA = 0, 1/2, 2/3, 3/4, and 1). The design is given in Table 1. For further discussion on the construction of component-amount designs via projections, see Prescott and Draper [16].
Table 1: Simplex-centroid design with four ingredients projected to a
component–amount design with three dimensions and five levels of AA
Simplex–centroid design with the proportions of 4 ingredients Component–amount design with the amounts of 3 ingredients
Run $x_{1}$ $x_{2}$ $x_{3}$ $x_{4}$ Run $a_{1}$ $a_{2}$ $a_{3}$ AA
1 1 0 0 0 1 1 0 0 1
2 0 1 0 0 2 0 1 0 1
3 0 0 1 0 3 0 0 1 1
4 0 0 0 1 4 0 0 0 0
5 1/2 1/2 0 0 5 1/2 1/2 0 1
6 1/2 0 0 1/2 6 1/2 0 0 1
7 1/2 0 1/2 0 7 1/2 0 1/2 1
8 0 1/2 1/2 0 8 0 1/2 1/2 1
9 0 1/2 0 1/2 9 0 1/2 0 1
10 0 0 1/2 1/2 10 0 0 1/2 1
11 1/3 1/3 1/3 0 11 1/3 1/3 1/3 1
12 1/3 1/3 0 1/3 12 1/3 1/3 0 2/3
13 1/3 0 1/3 1/3 13 1/3 0 1/3 2/3
14 0 1/3 1/3 1/3 14 0 1/3 1/3 2/3
15 1/4 1/4 1/4 1/4 15 1/4 1/4 1/4 3/4
@@ 2 OofA Experiments
The current literature on OofA experiments focuses primarily on the Pair-Wise Ordering model, first introduced by [18]. The Pair-Wise Ordering (PWO) model was formally named by Voelkel [19]. Let $\mathbf{c}=(c_{1},c_{2},\dots,c_{m})^{T}$ denote a permutation of the $m$ components. Let $\mathcal{P}$ be the set of all pairs $(j,k)$ where $1\leq j<k\leq m$.The PWO factor $z_{jk}(\mathbf{c})$ is defined as
[…]
@@
The parameter $\beta_{jk}$ shows how the pair-wise order of components $j$ and $k$ affect the response. For finding the optimal order Lin and Peng [20] discussed the topological sorting methods for the PWO model. Several research studies can be seen in the literature on the Optimality of PWO experimental designs. See, for instance, [21, 22].
@@ 2.1 The OofA mixture-amount experiments
An OofA mixture-amount experiment is a type of mixture experiment where we consider not only the effects of the component amounts on the response but also the effects of the order in which the components are added. The response surface would be a function of component amounts $a_{1},a_{2},\dots,a_{m}$ and the order of their addition in a mixture. We test the significant effect of the order of addition of components and the amounts of components on the response. Our concern in this study is how to incorporate OofA effects into the mixture-amount and component-amount models. Therefore, we construct the OofA mixture component-amount designs using mixture simplex designs. We use the modified PWO notations given by Rios and Lin [12]. Define (m1)(m-1) dimensional simplex $S$ and $\mathcal{P}$ be the set of all pairs $(j,k)$ where $1\leq j<k\leq m$. To get the full design matrix for fitting the mixture-amount model, we define PWO factor $z_{jk}(x,\mathbf{c})$ as
[…]
@@
[…]
@@ 2.2 Models for OofA mixture-amount experiments
We further constructed the model for the PWO and mixture component amounts. Without interaction between the mixture amount and PWO factor, the expected response additive model is developed as
@@
2.3 TA algorithm for OofA mixture-amount designs
Dueck and Scheuer [23] proposed a heuristic optimization algorithm called the TA algorithm. The modified algorithm for this study is given in Algorithm 2. It is widely applied to solving experimental design problems. This algorithm allows worse solutions to be accepted if only they fall within the threshold. Over time, the threshold is reduced, leading the algorithm toward a more refined solution (see Algorithm 1). Rios et al. [13] used TA algorithm to determine D-optimal OofA mixture designs. The D-optimality criterion maximizes the determinant of the information matrix. We use the TA algorithm to determine G-optimal OofA mixture-amount design with optimal number of runs $n\geq p$, where $p$ is the number of model parameters. The G-optimality criterion minimizes the maximum prediction variance over the design space.
$\text{G-optimality =}\quad\min_{\xi}\left(\max_{x\in\mathcal{X}}\mathbf{f}^{\top}(x)\left[\mathbf{F}^{\top}\mathbf{F}\right]^{-1}\mathbf{f}(x)\right)$
where $\xi$ denotes the experimental design, $\mathcal{X}$ is the design space, $\mathbf{f}(x)$ is the model vector evaluated at point $x$, and $\mathbf{F}$ represents the model matrix constructed from the selected design points.
Algorithm 1 Empirical Threshold Sequence Generation
1:Input: Number of iterations $n_{\text{iter}}$, full design $D_{\text{full}}$, design size $n$
2:Output: Empirical threshold sequence $\mathcal{T}^{*}$
3:Initialize $D_{0}\subset D_{\text{full}}$ such that $|D_{0}|=n$
4:Initialize empty set $\mathcal{T}\leftarrow\emptyset$
5:for $r=1$ to $n_{\text{iter}}$ do
6:  Generate the design $D_{1}$ by randomly exchanging $k$ rows in $D_{0}$ with $k$ rows from $D_{\text{full}}\setminus D_{0}$.
7:  Compute efficiency difference as $t_{r}=\left|\text{Efficiency}(D_{0})-\text{Efficiency}(D_{1})\right|$
8:  Update threshold set: $\mathcal{T}\leftarrow\mathcal{T}\cup\{t_{r}\}$
9:end for
10:Sort $\mathcal{T}$ in descending order: $\mathcal{T}_{\text{sorted}}$
11:Keep top 50% values $\mathcal{T}^{*}=\left\{t_{1},t_{2},\dots,t_{\lfloor n_{\text{iter}}/2\rfloor}\right\}$
12:Return $\mathcal{T}^{*}$
Algorithm 2 TA Algorithm for Finding G-Optimal OofA Mixture-Amount Design
1:Input: Full design $D_{\text{full}}$, number of design points $n$, number of iterations $n_{\text{iter}}$, model function $f(x)$, threshold sequence $\{t_{r}\}$
2:Output: Final design $D^{*}$ with improved G-efficiency
3:Randomly select $n$ rows from $D_{\text{full}}$ to form the initial design $D_{0}$
4:Construct the model matrix $M_{0}$ for $D_{0}$
5:Compute the inverse information matrix as $C_{0}=(M_{0}^{T}M_{0})^{-1}$
6:Compute the initial maximum prediction variance as $v_{\max}(D_{0})=\max_{x\in D_{\text{full}}}f(x)^{T}C_{0}f(x)$
7:for $r=1$ to $n_{\text{iter}}$ do
8:  Initialize: $D_{1}\leftarrow D_{0}$, $C_{1}\leftarrow C_{0}$, $v_{\max}(D_{1})\leftarrow v_{\max}(D_{0})$
9:  Generate a neighbor design $D_{1}\in N_{k}(D_{0})$ by randomly exchanging $k$ rows in $D_{0}$
10:  for each exchange $(x\in D_{0},y\in D_{\text{full}}\setminus D_{0})$ do
11:   Compute $v(y)=f(y)^{T}C_{0}f(y)$
12:   Update inverse matrix using Sherman–Morrison formula:
$C_{1}=C_{0}+C_{0}F_{1}(I_{2}+F_{2}^{T}C_{0}F_{1})^{-1}F_{2}^{T}C_{0}$
13:   Compute: $v_{\max}(D_{1})=\max_{x\in D_{\text{full}}}f(x)^{T}C_{1}f(x)$
14:  end for
15:  if $v_{\max}(D_{1})<v_{\max}(D_{0})+t_{r}$ then
16:   Accept $D_{1}$ as the new current design: $D_{0}\leftarrow D_{1}$, $C_{0}\leftarrow C_{1}$
17:  end if
18:end for
19:Return: Final design $D^{*}=D_{0}$ with associated G-efficiency